OpenStudy (anonymous):

Courtney can run 4.5 miles in 36 minutes .. How many minutes would it take her to run mini marathon of 12 miles at this same rate??

OpenStudy (lasttccasey):

Well the rate is distance over time so 4.5/36 miles per minute (mi/min). distance divided by rate will give you time. so D=12 R=.125mi/min T=? D/R=T Solve for T

OpenStudy (arabpride):

Well u have to start off by finding unit rate....

OpenStudy (arabpride):

divide 36 by 4.5...

OpenStudy (arabpride):

and u get 8...

OpenStudy (lasttccasey):

Rate is a measurement of something per unit time.

OpenStudy (anonymous):

4.5/36 = 12/x (4.5 miles to 36 min = 12 miles to x min) cross multiply (4.5)(x) = (12)(36) 4.5x = 432 x = 432/4.5 x = 96 minutes

OpenStudy (arabpride):

NOw u have to multiply 8 and 12, givvin u the number of minutes it would take to run 12 miles

OpenStudy (arabpride):

and.. 8 x 12 = 96

OpenStudy (arabpride):

so.....96 minutes to run 12 miles